OpenStudy (anonymous):

differentiate y=e^-2x ln (x^2)

OpenStudy (anonymous):

(e^(-2xln(x^2)))(-2ln(x^2)-4)

OpenStudy (anonymous):

\[y'(x)=\frac{2+\ln(x^2)}{e^2}\]

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

\[\frac{ 2e ^{-2x} }{ x }-2lnx ^{2}.e ^{-2x}\]